OpenStudy (anonymous):

2/3x+5=1/3x+14 Please go slow because I don't understand these.. :(

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

the trick to these is to get all your variables on one side first \[2/(3x) + 5 =1/(3x) + 14\] subtract 5 from both sides \[2/(3x) =1/(3x) + 9\] now subtract 1/(3x) on each side. \[1/(3x) = 9\] multiply by 3x on both sides \[1 = 27x\] divide by 27 \[1/27 = x\]

OpenStudy (anonymous):

No you subtract a number with an x next to it.

OpenStudy (anonymous):

 is your problem the first line or second?